# PRISM: Prompt-Refined In-Context System Modelling for Financial Retrieval

### v1

@@ A.6.1 Descriptive Statistics
All configurations exhibit low standard deviations ($s<0.011$) and coefficients of variation below 1.6%, indicating highly stable and reproducible results across runs. The 95% confidence intervals are narrow, confirming that the estimated means are statistically reliable. Run 12 achieves the lowest mean performance ($\bar{x}=0.68005$), while Runs 15–19 yield consistently higher means (0.70246–0.71163). The non-overlapping confidence intervals between Run 12 and the later runs demonstrate that the configuration changes introduced after Run 12 resulted in a statistically meaningful performance gain. Subsequent runs (15–19) show only minor numerical variation, suggesting that performance has stabilized and that the system’s behaviour is reproducible.
A.6.2 Statistical Significance
@@
where $\bar{x}_{1},s_{1},n_{1}$ and $\bar{x}_{2},s_{2},n_{2}$ denote the mean, standard deviation, and sample size of each group. The degrees of freedom are approximated using the Welch–Satterthwaite equation:
@@
A.6.3 Interpretation of Statistical Significance
@@
A.7 Production Readiness Analysis
This section evaluates the production readiness of the PRISM components across prompt engineering, ICL retrieval, and multi-agent system modelling.
A.7.1 Prompt Engineering
@@
A.7.2 In-Context Learning Sample Retrieval
@@
A.7.3 Multi-Agent System (MAS) Modelling
While MAS offers methodological advantages, it is not yet operationally viable for large-scale deployment. Its coordination overhead, sensitivity to prompt–architecture interactions, and variable latency make it difficult to guarantee predictable performance and cost efficiency. These limitations pose challenges for enterprise-grade integration, suggesting that MAS workflows remain better suited to research and prototyping rather than near-term production use.
A.7.4 Toward Production Readiness
Several directions can help bridge the gap between experimental performance and production maturity. Firstly, automated or data-driven prompt optimization (e.g., search-based or Reinforcement Learning-assisted refinement) could reduce manual engineering effort while ensuring consistency across model updates. Secondly, for ICL sample retrieval, future work can focus on developing adaptive exemplar selection mechanisms that dynamically adjust the number and relevance of exemplars based on more detailed requirements such as the query complexity and token constraints, tailoring retrieval depth to query characteristics. Thirdly, for MAS modelling, progress in agent coordination frameworks, caching mechanisms, and lightweight state management could reduce overhead and latency, making the workflows more predictable and cost-efficient. Integrating these components into a unified monitoring and evaluation framework would also enhance reliability and traceability in real-world deployments. In summary, these improvements would strengthen system scalability and operational control, moving PRISM closer to production-grade readiness.
A.8 Prompt Template
The non-agentic prompt templates are designed for single LLM workflows, where the model independently performs the complete ranking process using structured system instructions. These prompts emphasize logical consistency, domain alignment and interpretable reasoning, progressing from ReAct-style instructions to more structured CoT and ToT variants that balance interpretability and precision. Each iteration refines clarity, output constraints, and contextual cues to improve both document and chunk ranking performance. In contrast, the agentic prompt templates extend this scaffolding to a multi-role architecture, where specialized agents inherit the underlying reasoning framework while contributing from distinct role and domain-specific perspectives such as Financial Analyst, Risk Analyst, or Evidence Extractor. This modular design facilitates more interpretable and coordinated reasoning across complex financial scenarios but remains sensitive to prompt compatibility and inter-agent communication constraints.
@@ A.8 Prompt Template
$P_{3}$ Task: - Internally reason about which documents are most relevant, using explicit chain-of-thought reasoning, Tree-of-Thought simulation, and ReAct (Reason+Act) strategy to break down and analyze complex financial queries and ranking tasks. - Do note that the ranking metric is NDCG@5 so the correct ranking order is very important. - Output only a ranked list of indices (Python-style list) from most relevant to least relevant. - Do not output explanations, scores, or JSON mappings. Domain ranking hints: - Use 10-K for strategy, risks, or full-year data. - Use 10-Q for recent quarterly updates. - Use 8-K for specific events/announcements. - Use DEF14A for governance, board, and compensation. - Use Earnings for quarterly results or management commentary. Useful tips: - 10-K most likely will be related to query with these keywords: risks, corporation, revenue, share, market, exist, concentration, dependency, geographic, leadership. - 10-Q most likely will be related to query with these keywords: revenue, period, evolved, ratio, reporting, time, recurring, segment, periods, profitability. - 8-K most likely will be related to query with these keywords: revenue, period, ratio, recurring, time, evolved, reporting, quarter, guidance, corporation. - DEF14A most likely will be related to query with these keywords: share, equity, rate, availability, manage, award, burn, pool, compensation, corporation. - Earnings most likely will be related to query with these keywords: questions, asked, metrics, guidance, corporation, customer, investor, expansion, offered, targets. Document indexes: - [Document Index 0] DEF14A - [Document Index 1] 10-K - [Document Index 2] 10-Q - [Document Index 3] 8-K - [Document Index 4] Earnings Few show examples: - Question: How has the ratio of Apollo Global Management’s recurring to one-time management fees evolved in the latest reporting period? | Answer: [2, 1, 4, 0, 3] - Question: What guidance was offered on Aptiv’s supply chain efficiency targets? | Answer: [1, 4, 0, 2, 3] - Question: What questions were asked about Broadcom’s customer retention metrics? | Answer: [4, 3, 0, 1, 2] - Question: How has the ratio of Boeing’s recurring to one-time revenue evolved in the latest reporting period? | Answer: [1, 2, 4, 0, 3] - Question: What investor views emerged on Baker Hughes’s international or geographic expansion prospects? | Answer: [4, 1, 2, 3, 0] Persistence - You are an agent - please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Only terminate your turn when you are sure that the problem is solved. - Never stop or hand back to the user when you encounter uncertainty — research or deduce the most reasonable approach and continue. - Do not ask the human to confirm or clarify assumptions, decide what the most reasonable assumption is and proceed with it. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{4}$ Task: - Internally reason about which documents are most relevant, using explicit chain-of-thought reasoning, Tree-of-Thought simulation, and ReAct (Reason+Act) strategy to break down and analyze complex financial queries and ranking tasks. - Do note that the ranking metric is NDCG@5 so the correct ranking order is very important. - Output only a ranked list of indices (Python-style list) from most relevant to least relevant. - Do not output explanations, scores, or JSON mappings. Domain ranking hints: - Use 10-K for strategy, risks, or full-year data. - Use 10-Q for recent quarterly updates. - Use 8-K for specific events/announcements. - Use DEF14A for governance, board, and compensation. - Use Earnings for quarterly results or management commentary. Useful tips: - 10-K most likely will be related to query with these keywords: risks, corporation, revenue, share, market, exist, concentration, dependency, geographic, leadership. - 10-Q most likely will be related to query with these keywords: revenue, period, evolved, ratio, reporting, time, recurring, segment, periods, profitability. - 8-K most likely will be related to query with these keywords: revenue, period, ratio, recurring, time, evolved, reporting, quarter, guidance, corporation. - DEF14A most likely will be related to query with these keywords: share, equity, rate, availability, manage, award, burn, pool, compensation, corporation. - Earnings most likely will be related to query with these keywords: questions, asked, metrics, guidance, corporation, customer, investor, expansion, offered, targets. Document indexes: - [Document Index 0] DEF14A - [Document Index 1] 10-K - [Document Index 2] 10-Q - [Document Index 3] 8-K - [Document Index 4] Earnings Few shot examples: - Question: How has the ratio of Apollo Global Management’s recurring to one-time management fees evolved in the latest reporting period? | Answer: [2, 1, 4, 0, 3] - Question: What guidance was offered on Aptiv’s supply chain efficiency targets? | Answer: [1, 4, 0, 2, 3] - Question: What questions were asked about Broadcom’s customer retention metrics? | Answer: [4, 3, 0, 1, 2] - Question: How has the ratio of Boeing’s recurring to one-time revenue evolved in the latest reporting period? | Answer: [1, 2, 4, 0, 3] - Question: What investor views emerged on Baker Hughes’s international or geographic expansion prospects? | Answer: [4, 1, 2, 3, 0] Important: - Please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. - Your accurate ranking of the chunks are the most important task. Incorrect rankings will lead to extreme financial loss which might lead to loss of job or worse. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{2}$ Task: - Input: A question and a list of text chunks with indices. - Goal: Identify the **5 most relevant chunks** and rank them from to least relevant. - Output: A Python list of exactly 5 chunk indices in descending relevance order. Constraints: - Always return exactly 5 indices. - Output must be a flat Python list of integers (e.g., ‘[7, 12, 3, 9, 14]‘). - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? | Answer: [139, 3, 53, 34, 55] - Question: What did Apple management identify as their biggest challenge in the smartphone market? | Answer: [2, 11, 5, 9, 14] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? | Answer: [6, 1, 13, 7, 10] Persistence - You are an agent - please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Only terminate your turn when you are sure that the problem is solved. - Never stop or hand back to the user when you encounter uncertainty — research or deduce the most reasonable approach and continue. - Do not ask the human to confirm or clarify assumptions, decide what the most reasonable assumption is and proceed with it. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{3}$ Task: - Input A question and a list of text chunks with indices. - Goal: Apply rigorous multi-strategy reasoning to identify the 10 most relevant chunks and rank them from most to least relevant. - Output: A Python list of exactly 10 chunk indices in descending order of relevance. Constraints: - Always return exactly 10 indices. - Output must be a flat Python list of integers. - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? | Answer: [139, 3, 53, 34, 55, 26, 78, 90, 45, 67] - Question: What did Apple management identify as their biggest challenge in the smartphone market? | Answer: [2, 11, 5, 9, 14, 1, 4, 7, 3, 8] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? | Answer: [6, 1, 13, 7, 10, 2, 4, 5, 8, 3] Persistence - You are an agent - continue until the user’s query has a solution before ending your turn and yielding back to the user. - Only terminate your turn once you are sure the problem is solved. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3, 5, 6, 7, 8, 9] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{4}$ Task: - Input A question and a list of text chunks with indices. - Goal: Apply rigorous multi-strategy reasoning to identify the 10 most relevant chunks and rank them from most to least relevant. - Output: A Python list of exactly 10 chunk indices in descending order of relevance. Constraints: - Always return exactly 10 indices. - Output must be a flat Python list of integers. - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? | Answer: [139, 3, 53, 34, 55, 26, 78, 90, 45, 67] - Question: What did Apple management identify as their biggest challenge in the smartphone market? | Answer: [2, 11, 5, 9, 14, 1, 4, 7, 3, 8] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? | Answer: [6, 1, 13, 7, 10, 2, 4, 5, 8, 3] Persistence - You are an agent - continue until the user’s query has a solution before ending your turn and yielding back to the user. - Only terminate your turn once you are sure the problem is solved. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3, 5, 6, 7, 8, 9] Do not return any text outside of the list. Any other format is incorrect.

### v2

@@ A.11.1 Descriptive Statistics
All configurations exhibit low standard deviations ($s<0.011$) and coefficients of variation below 1.6%, indicating highly stable and reproducible results across runs. The 95% confidence intervals are narrow, confirming that the estimated means are statistically reliable. Run 12 achieves the lowest mean performance ($\bar{x}=0.68005$), while Runs 15–19 yield consistently higher means (0.70246–0.71163). The non-overlapping confidence intervals between Run 12 and the later runs demonstrate that the configuration changes introduced after Run 12 resulted in a statistically meaningful performance gain. Subsequent runs (15–19) show only minor numerical variation, suggesting that performance has stabilized and that the system’s behavior is reproducible.
A.11.2 Statistical Significance
@@
where $\bar{x}_{1},s_{1},n_{1}$ and $\bar{x}_{2},s_{2},n_{2}$ denote the mean, standard deviation, and sample size of each group. The degrees of freedom are approximated using the Welch-Satterthwaite equation:
@@
A.11.3 Interpretation of Statistical Significance
@@
A.12 Production Readiness Analysis
This section evaluates the production readiness of the PRISM components across prompt engineering, ICL retrieval, and multi-agent system modeling. While our best-performing configuration relies primarily on prompt engineering and selective ICL, we include all three components to provide a comprehensive empirical analysis of the design space. This approach enables practitioners to make informed decisions about which components to deploy based on their specific constraints and requirements.
A.12.1 Prompt Engineering
@@
A.12.2 In-Context Learning Sample Retrieval
@@
A.12.3 Non-Agentic Workflow
This workflow represents the most production-ready configuration within PRISM, leveraging prompt engineering and selective ICL retrieval in a single-LLM architecture. The workflow’s production readiness stems from three key advantages. First, it inherits prompt engineering’s operational maturity that requires no architectural changes, delivering predictable latency, and adapting seamlessly to new foundation models. Second, it incorporates ICL retrieval selectively, allowing organizations to activate exemplar-based guidance only for complex queries where the performance gains justify additional token costs, maintaining cost efficiency for routine tasks. Third, the single-agent design ensures more deterministic behavior with minimal variance, critical for enterprise deployments requiring consistent service-level agreements.
Our four prompt variants ($P_{1}$$P_{4}$) demonstrate progressive sophistication while maintaining production viability. The modular prompt design enables flexible production deployment strategies where organizations can start with simpler prompts ($P_{1}$) for cost-sensitive applications, adopt intermediate complexity ($P_{2}$, $P_{3}$) for balanced performance-cost trade-offs, or deploy advanced reasoning ($P_{4}$) for high-stakes queries. Combined with adaptive ICL activation varying exemplar count based on query complexity or token budgets, the non-agentic workflow provides a scalable, operationally reliable foundation suitable for immediate enterprise integration. We also note that prompt length management is critical as discussed in Appendix A.13.5, where LLM performance can degrade with increasing context length, reinforcing the design rationale for $P_{4}$’s streamlined structure over more verbose alternatives.
A.12.4 Agentic Workflow
The agentic workflow extends the non-agentic prompt scaffolding to a multi-role architecture, where specialized agents inherit the underlying ReAct reasoning framework while contributing distinct domain-specific perspectives. Each agent scores chunks on a 1–10 scale based on its specialized focus area. This modular design facilitates interpretable and coordinated reasoning across complex financial scenarios, enabling role-based specialization that mirrors real-world expert collaboration. Full agent definitions and detailed workflow architectures are provided in Section A.7.
Our systematic evaluation indicates that MAS is not yet operationally viable for large-scale deployment. Coordination overhead, sensitivity to prompt-architecture interactions, and variable latency make it difficult to guarantee predictable performance and cost efficiency. In particular, error propagation in deep pipelines, inconsistent scoring with smaller models, and diminishing returns from agent specialization in chunk ranking all pose challenges for enterprise-grade integration. A detailed analysis of these failure modes and their implications is provided in Section A.13.
A.13 Limitations
A.13.1 Data Preprocessing
A.13.2 Non-Agentic Workflow
Although the non-agentic configuration achieved the strongest overall performance, retrieval fidelity remains a notable limitation. Our best ICL setup relies heavily on the quality of the TE3-S model, and currently uses a simple L2 distance-based embedding lookup to retrieve relevant chunks. More advanced retrieval strategies could yield higher-quality candidates. A future improvement could involve hybrid retrieval that integrates dense semantic embeddings with lexical methods such as BM25 (Robertson et al., 1995), combined through Reciprocal Rank Fusion (Cormack et al., 2009). This would provide a more balanced retrieval signal, capturing both semantic context and domain-specific financial terminology that purely dense methods may overlook.
A.13.3 Prompt-Architecture Asymmetry
A known limitation in our experimental design is that agentic workflow experiments consistently use $P_{1}$, while the best non-agentic result uses $P_{4}$. As discussed in Appendix A.8, this was not an arbitrary choice: applying $P_{4}$ in multi-agent settings caused token limit violations and degraded inter-agent coordination, as evidenced by Run 32. Running additional controlled experiments to fully resolve this asymmetry was not feasible, both due to the practical constraints described above and because the FinAgentBench evaluation environment is no longer available (see §A.13, Benchmark Reproducibility). Cross-workflow comparisons should therefore be read as reflecting how prompt complexity interacts with agent architecture, rather than as a head-to-head evaluation under matched conditions.
A.13.4 Agentic Workflow
Our agentic experiments revealed several critical challenges and failure modes. First, multi-agent systems are highly sensitive to prompt-architecture compatibility: the $P_{4}$ prompt, which consistently improved non-agentic performance, led to substantial performance drops in MAS settings, suggesting that constraint-heavy prompts can disrupt inter-agent communication. Second, $A_{2}$’s three-stage deep pipeline exhibited catastrophic recall failures through error propagation, where filtering agents rejected relevant chunks that downstream scoring agents never observed, particularly when filtering inconsistency compounded across stages. Third, coordination overhead with smaller models proved significant, as GPT-4o-mini struggled with multi-agent coordination across all architectures, often producing inconsistent scores that degraded when averaged, leading single-agent baselines to frequently outperform complex MAS configurations. Finally, over-engineering in chunk ranking emerged as a consistent pattern, where simpler architectures like $A_{4}$ matched or exceeded deeper variants ($A_{2}$, $A_{3}$) across most metrics, suggesting that agent specialization benefits are outweighed by coordination costs for fine-grained tasks.
Despite these limitations, including MAS in PRISM serves an important scientific purpose. It empirically identifies the boundary conditions under which multi-agent reasoning provides value versus introducing failure modes. Our systematic evaluation reveals that document-level ranking consistently benefits from MAS, particularly with larger models (GPT-5), where coarse-grained tasks leverage role specialization effectively. In contrast, chunk-level ranking suffers from coordination overhead and error propagation, especially with smaller models or deeper graph topologies. These findings inform our recommendation to reserve MAS for document-level or coarse-grained ranking tasks with capable models (GPT-5), while preferring simpler single-agent or minimal two-agent configurations ($A_{4}$) for fine-grained chunk ranking, especially when computational budgets or model capabilities are limited. Future work could explore automated prompt-architecture optimization techniques to jointly tune prompt complexity and agent topology, as well as developing standardized inter-agent communication protocols that remain robust across varying prompt constraints and support adaptive coordination mechanisms tailored for distributed reasoning.
A.13.5 Context Rot with Complicated Prompts
Recent research on context rot (Hong et al., 2025) demonstrates that LLM performance degrades as input length increases, even on simple tasks. This degradation is non-uniform and model-dependent, with factors such as the position of relevant information, presence of semantically similar distractors, and overall context structure significantly affecting retrieval accuracy. These findings have direct implications for PRISM’s design choices.
Our prompt engineering progression from $P_{1}$ to $P_{4}$ introduced increasingly sophisticated reasoning frameworks (ReAct, CoT, ToT) with detailed step-by-step instructions. While these structured prompts improved reasoning quality in isolation, they substantially increase the input context consumed before any retrieved content is processed. Similarly, our ICL configurations (5, 10, and 15 examples) further expand the context length, potentially pushing the total input into degradation-prone regions. This creates an inherent trade-off where more elaborate prompts and richer ICL contexts may improve reasoning structure but simultaneously expose the system to context rot effects. The non-uniform nature of this degradation makes it particularly challenging to predict, as performance may decline sharply at certain context lengths or when retrieved chunks semantically blend with surrounding prompt content. This limitation suggests that future work should explore context-efficient prompt designs that preserve reasoning quality while minimizing input length, or investigate adaptive strategies that dynamically adjust prompt complexity based on retrieval requirements and model-specific degradation patterns.
A.13.6 Model Provider Generalization
All experiments were conducted exclusively using OpenAI models due to infrastructure constraints (Azure ecosystem availability and limited computational credits). While OpenAI models are well-benchmarked on financial tasks and provide a stable experimental baseline, this limits the scientific generality of our findings. In particular, behaviors such as model-size effects on MAS coordination and ICL sensitivity are likely model-dependent, as they rely on instruction-following capacity that may differ substantially across architectures. Performance characteristics may therefore differ with open-source alternatives such as LLaMA or Mistral, which have different architectural properties and training data distributions. The training-free nature of PRISM facilitates straightforward adaptation to other providers, and we encourage future work to validate our findings across diverse model families to establish broader applicability.
A.13.7 Prompt Template Specificity
Several prompt templates include concrete illustrative examples drawn from each of the training set of experimented datasets, such as company-specific few-shot demonstrations with entity names and index labels. While these examples were selected to reflect the distributional properties of the training data rather than to target specific test queries, there remains a risk that such examples are better calibrated to the evaluation distribution than to other financial IR tasks. Practitioners applying PRISM to different datasets should verify that few-shot examples are sampled from a held-out exemplar pool representative of their target distribution.
A.13.8 Benchmark Reproducibility
The FinAgentBench evaluation environment became unavailable following the competition’s conclusion, preventing further experiments or independent replication of reported scores. Ground truth labels for the public and private test sets were not disclosed by the benchmark organizers. As a result, the competition-phase results (Table 9 and Run 19 in Table 3) cannot be independently verified or extended without access to the original evaluation system. Our ablation experiments on the validation split remain reproducible given the released code.
A.13.9 Architectural Design
Architecturally, our results indicate a saturation bottleneck in the current MAS design. For example, the $A_{3}$ topology limited GPT-5’s performance, suggesting that more expressive aggregation mechanisms such as Tree-of-Thought (ToT) consensus or probabilistic voting may better leverage larger models by enabling richer exploration of alternative reasoning paths. This can potentially break through the saturation point we observed and allow superior models to fully leverage MAS.
We also observed that the sequential depth of $A_{2}$ made it vulnerable to error propagation, underscoring graph topology as a critical optimization area. Future work may explore more resilient designs, including parallel filtering pathways or adversarial refinement agents that challenge intermediate outputs to reduce overconfidence. Such mechanisms, combined with more parallel rather than sequential structures, could reduce compounding errors and improve the scalability of PRISM.
A.14 Prompt Template
@@
$P_{3}$ Task: - Internally reason about which documents are most relevant, using explicit chain-of-thought reasoning, Tree-of-Thought simulation, and ReAct (Reason+Act) strategy to break down and analyze complex financial queries and ranking tasks. - Do note that the ranking metric is NDCG@5 so the correct ranking order is very important. - Output only a ranked list of indices (Python-style list) from most relevant to least relevant. - Do not output explanations, scores, or JSON mappings. Domain ranking hints: - Use 10-K for strategy, risks, or full-year data. - Use 10-Q for recent quarterly updates. - Use 8-K for specific events/announcements. - Use DEF14A for governance, board, and compensation. - Use Earnings for quarterly results or management commentary. Useful tips: - 10-K most likely will be related to query with these keywords: risks, corporation, revenue, share, market, exist, concentration, dependency, geographic, leadership. - 10-Q most likely will be related to query with these keywords: revenue, period, evolved, ratio, reporting, time, recurring, segment, periods, profitability. - 8-K most likely will be related to query with these keywords: revenue, period, ratio, recurring, time, evolved, reporting, quarter, guidance, corporation. - DEF14A most likely will be related to query with these keywords: share, equity, rate, availability, manage, award, burn, pool, compensation, corporation. - Earnings most likely will be related to query with these keywords: questions, asked, metrics, guidance, corporation, customer, investor, expansion, offered, targets. Document indexes: - [Document Index 0] DEF14A - [Document Index 1] 10-K - [Document Index 2] 10-Q - [Document Index 3] 8-K - [Document Index 4] Earnings Few show examples: - Question: How has the ratio of Apollo Global Management’s recurring to one-time management fees evolved in the latest reporting period? — Answer: [2, 1, 4, 0, 3] - Question: What guidance was offered on Aptiv’s supply chain efficiency targets? — Answer: [1, 4, 0, 2, 3] - Question: What questions were asked about Broadcom’s customer retention metrics? — Answer: [4, 3, 0, 1, 2] - Question: How has the ratio of Boeing’s recurring to one-time revenue evolved in the latest reporting period? — Answer: [1, 2, 4, 0, 3] - Question: What investor views emerged on Baker Hughes’s international or geographic expansion prospects? — Answer: [4, 1, 2, 3, 0] Persistence - You are an agent - please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Only terminate your turn when you are sure that the problem is solved. - Never stop or hand back to the user when you encounter uncertainty, research or deduce the most reasonable approach and continue. - Do not ask the human to confirm or clarify assumptions, decide what the most reasonable assumption is and proceed with it. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{4}$ Task: - Internally reason about which documents are most relevant, using explicit chain-of-thought reasoning, Tree-of-Thought simulation, and ReAct (Reason+Act) strategy to break down and analyze complex financial queries and ranking tasks. - Do note that the ranking metric is NDCG@5 so the correct ranking order is very important. - Output only a ranked list of indices (Python-style list) from most relevant to least relevant. - Do not output explanations, scores, or JSON mappings. Domain ranking hints: - Use 10-K for strategy, risks, or full-year data. - Use 10-Q for recent quarterly updates. - Use 8-K for specific events/announcements. - Use DEF14A for governance, board, and compensation. - Use Earnings for quarterly results or management commentary. Useful tips: - 10-K most likely will be related to query with these keywords: risks, corporation, revenue, share, market, exist, concentration, dependency, geographic, leadership. - 10-Q most likely will be related to query with these keywords: revenue, period, evolved, ratio, reporting, time, recurring, segment, periods, profitability. - 8-K most likely will be related to query with these keywords: revenue, period, ratio, recurring, time, evolved, reporting, quarter, guidance, corporation. - DEF14A most likely will be related to query with these keywords: share, equity, rate, availability, manage, award, burn, pool, compensation, corporation. - Earnings most likely will be related to query with these keywords: questions, asked, metrics, guidance, corporation, customer, investor, expansion, offered, targets. Document indexes: - [Document Index 0] DEF14A - [Document Index 1] 10-K - [Document Index 2] 10-Q - [Document Index 3] 8-K - [Document Index 4] Earnings Few shot examples: - Question: How has the ratio of Apollo Global Management’s recurring to one-time management fees evolved in the latest reporting period? — Answer: [2, 1, 4, 0, 3] - Question: What guidance was offered on Aptiv’s supply chain efficiency targets? — Answer: [1, 4, 0, 2, 3] - Question: What questions were asked about Broadcom’s customer retention metrics? — Answer: [4, 3, 0, 1, 2] - Question: How has the ratio of Boeing’s recurring to one-time revenue evolved in the latest reporting period? — Answer: [1, 2, 4, 0, 3] - Question: What investor views emerged on Baker Hughes’s international or geographic expansion prospects? — Answer: [4, 1, 2, 3, 0] Important: - Please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. - Your accurate ranking of the chunks are the most important task. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{2}$ Task: - Input: A question and a list of text chunks with indices. - Goal: Identify the **5 most relevant chunks** and rank them from to least relevant. - Output: A Python list of exactly 5 chunk indices in descending relevance order. Constraints: - Always return exactly 5 indices. - Output must be a flat Python list of integers (e.g., ‘[7, 12, 3, 9, 14]‘). - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? — Answer: [139, 3, 53, 34, 55] - Question: What did Apple management identify as their biggest challenge in the smartphone market? — Answer: [2, 11, 5, 9, 14] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? — Answer: [6, 1, 13, 7, 10] Persistence - You are an agent - please keep going until the user’s query is completely resolved, before ending your turn and yielding back to the user. - Only terminate your turn when you are sure that the problem is solved. - Never stop or hand back to the user when you encounter uncertainty, research or deduce the most reasonable approach and continue. - Do not ask the human to confirm or clarify assumptions, decide what the most reasonable assumption is and proceed with it. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{3}$ Task: - Input A question and a list of text chunks with indices. - Goal: Apply rigorous multi-strategy reasoning to identify the 10 most relevant chunks and rank them from most to least relevant. - Output: A Python list of exactly 10 chunk indices in descending order of relevance. Constraints: - Always return exactly 10 indices. - Output must be a flat Python list of integers. - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? — Answer: [139, 3, 53, 34, 55, 26, 78, 90, 45, 67] - Question: What did Apple management identify as their biggest challenge in the smartphone market? — Answer: [2, 11, 5, 9, 14, 1, 4, 7, 3, 8] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? — Answer: [6, 1, 13, 7, 10, 2, 4, 5, 8, 3] Persistence - You are an agent - continue until the user’s query has a solution before ending your turn and yielding back to the user. - Only terminate your turn once you are sure the problem is solved. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3, 5, 6, 7, 8, 9] Do not return any text outside of the list. Any other format is incorrect.
@@
$P_{4}$ Task: - Input A question and a list of text chunks with indices. - Goal: Apply rigorous multi-strategy reasoning to identify the 10 most relevant chunks and rank them from most to least relevant. - Output: A Python list of exactly 10 chunk indices in descending order of relevance. Constraints: - Always return exactly 10 indices. - Output must be a flat Python list of integers. - No explanations, no text, no scores, no JSON. Any non-list output is incorrect. Few-Shot Examples: - Question: What investor views emerged on Arthur J. Gallagher & Co.’s international expansion prospects? — Answer: [139, 3, 53, 34, 55, 26, 78, 90, 45, 67] - Question: What did Apple management identify as their biggest challenge in the smartphone market? — Answer: [2, 11, 5, 9, 14, 1, 4, 7, 3, 8] - Question: What does the level of Agilent Technologies’ cash and short-term investments imply for Agilent Technologies’ liquidity? — Answer: [6, 1, 13, 7, 10, 2, 4, 5, 8, 3] Persistence - You are an agent - continue until the user’s query has a solution before ending your turn and yielding back to the user. - Only terminate your turn once you are sure the problem is solved. - Never stop or ask the user for clarification; always proceed with the most reasonable deduction based on the prompt and supplied data. Strict Output Format: Return only a Python-style list of indices. Example: [1, 0, 2, 4, 3, 5, 6, 7, 8, 9] Do not return any text outside of the list. Any other format is incorrect.